# Harmony: A Joint Self-Supervised and Weakly-Supervised Framework for Learning General Purpose Visual Representations

### v1

@@ Abstract
Vision-language contrastive learning frameworks like CLIP enable learning representations from natural language supervision, and provide strong zero-shot classification capabilities. However, due to the nature of the supervisory signal in these paradigms, they lack the ability to learn localized features, leading to degraded performance on dense prediction tasks like segmentation and detection. On the other hand, self-supervised learning methods have shown the ability to learn granular representations, complementing the high-level features in vision-language training. In this work, we present Harmony, a framework that combines vision-language training with discriminative and generative self-supervision to learn visual features that can be generalized across vision downstream tasks. Our framework is specifically designed to work on web-scraped data by not relying on negative examples and addressing the one-to-one correspondence issue using soft CLIP targets generated by an EMA model. We comprehensively evaluate Harmony across various vision downstream tasks and find that it significantly outperforms the baseline CLIP and the previously leading joint self and weakly-supervised methods, MaskCLIP and SLIP. Specifically, when comparing against these methods, Harmony shows superior performance in fine-tuning and zero-shot classification on ImageNet-1k, semantic segmentation on ADE20K, and both object detection and instance segmentation on MS-COCO, when pre-training a ViT-S/16 on CC3M. We also show that Harmony outperforms other self-supervised learning methods like iBOT and MAE across all tasks evaluated. Our code is publicly available at https://github.com/MohammedSB/Harmony.
Figure 1: Harmony compared to previous methods. We show our Harmony approach next to previously leading joint methods MaskCLIP [17] and SLIP [37]. Harmony optimizes five different objectives simultaneously, outperforming the other approaches across all vision downstream tasks.
@@ 1 Introduction
Self-supervised and weakly-supervised pre-training have recently shown remarkable success at learning visual representations without direct supervision [40, 38, 6, 9, 10, 23, 20, 55, 24, 37, 17]. As vision training datasets continue to scale, it becomes progressively more difficult and expensive to provide manual supervision in the form of labels, making the development of robust self-supervised learning (SSL) and weakly-supervised learning (WSL) techniques more integral.
@@
Weakly-supervised learning, or text-guided learning, was popularized by CLIP [40] and learns visual and textual representations using contrastive loss by maximizing the similarity between image-captions pairs and minimizing the similarity of non-paired image-captions [40, 12]. Because this approach relies on semantic captions as supervisory signal, weakly-supervised models are strong at high-level tasks like image classification, but significantly underperform on dense, low-level prediction tasks that require localized features [40, 47, 48]. In other words, these paradigms are good at learning what objects are present in a visual input, but not where they are. One approach of introducing local information into WSL frameworks is to combine it with self-supervised learning [17, 37, 53]. Unlike weakly-supervised learning that maps across modalities (e.g. image to text), SSL maps to the same visual modality, making it more granular and localized for visual tasks [6, 24, 2].
@@
Recent works in self-supervised learning formulate the pre-training task as either discriminative or generative [39, 16]. For discriminative SSL methods, the model learns visual representations from images by differentiating between positive images pairs, and optionally repelling negative pairs [9, 20, 6, 38, 23, 10]. On the other hand, generative approaches learn visual representations by masking certain parts of an image and learning to reconstruct the missing parts, given the original image [24, 44, 21]. Because SSL methods do not require manually annotated labels, they can be used for training large neural networks on huge image datasets scraped from the web [38].
@@
Combining weakly-supervised and self-supervised learning has led to better performing learning paradigms [17, 37, 53]. For example, SLIP [37] combines CLIP with SimCLR, leading to significantly better zero-shot capabilities, and overall more accurate models across multiple downstream tasks. More recently, MaskCLIP combines CLIP with masked self-distillation, leading to further improvements in a variety of downstream evaluations.
@@
Building on the aforementioned works, we present a novel framework, which we will call Harmony, that combines discriminative and generative self-supervision with weakly-supervised learning in order to learn general purpose visual representations from image-captions pairs gathered from the web. Our approach is designed to learn semantic visual representations useful for high-level vision tasks and fine-grained, low-level representation for dense prediction tasks at the same time.
@@
Previous approaches like MaskCLIP [17] and SLIP [37] rely on one-to-one correspondences and hard negative examples which fail to model the inherent semantic relationship between non-paired samples [1]. This is because for a certain image, captions of other images in the batch could still describe this image with varying degrees, especially when the batch size is large. Moreover, particularly in an uncurated data setting, the caption paired with a certain image could simply be incorrect, or the description could be only loosely associated with the image. To remedy these issues, we designed our framework to utilize one-to-many relationships by incorporating soft CLIP targets, and to not rely on negative examples by using self-distillation methods like iBOT [55] rather than SimCLR [9] like in SLIP [37]. We argue that this makes Harmony better designed for web-scraped data like CC3M [43].
@@
  • We present Harmony, a joint self-supervised and weakly-supervised framework that learns global and local features, generalizing across vision tasks including classification, segmentation, and detection.
  • Harmony is built not to use negative examples and to rely less on one-to-one correspondence in the text-guided contrastive path, leading to outperforming previous best methods SLIP and MaskCLIP when pre-training on the web-scraped CC3M [43] dataset.
@@ 2 Related Works
Discriminative self-supervision. Recent discriminative SSL approaches learn visual representations by discriminating between images that are positive pairs of each other, usually defined as different augmentations of the same image, and optionally pushing away negative pairs. [20, 6, 23, 10, 9]. Methods like SimCLR [9] and MoCo [23, 10] use contrastive learning to maximize the embedding similarity between positive pairs while simultaneously reducing the similarity of negative pairs. These approaches ignore the inherent similarity between different images and treat the discriminative problem as a one-to-one scheme, where an image is only similar to it’s positive pair. Other approaches like BYOL [20] and DINO [6] use self-distillation to only maximize agreement between positive pairs, avoiding that issue. These methods use a momentum encoder, usually refereed to as a target or teacher encoder, to generate embedding targets for an online or student encoder that is trained simultaneously with the teacher encoder. The learning task in these paradigms is to maximize the similarity between embeddings from both the teacher and student encoders given positive image pairs. Self-distillation methods regularly outperform contrastive methods, and reach accuracies that are competitive with supervised learning [6, 55, 20].
@@
Generative self-supervision. Rather than discriminating between images, generative SSL methods learn visual representations by masking certain parts of an images (or certain images in a video) and then learn to reconstruct the masked portions given the original signal [24, 44, 21]. This reconstruction task can be viewed as a proxy for learning about visual features from images, improving representation ability and, therefore, performance on downstream tasks. A common approach in this paradigm is called Masked Autoencoder (MAE) [24], which utilizes ViT’s patching paradigm to mask a certain percentage of patches (usually 75%) and minimizes the L2 loss on predicted pixels from the masked patches. Generative SSL approaches are highly scalable due to their lightweight memory usage compared to discriminative approaches, and reach fine-tuning accuracies rivaling those methods [24].
@@
Joint self-supervision methods. Many previous works have combined discriminative and generative self-supervised learning objectives, resulting in improved performance on downstream classification and semantic segmentation tasks [11, 27, 36]. CMAE [27] and CAE [11] combine MAE’s pixel reconstruction task with contrastive learning, outperforming either method in isolation. CAN [36] further adds a noise prediction task and masks both views in the contrastive learning objective. These work highlight the fact that SSL paradigms can be complimentary to each other.
@@
Weak and self-supervised learning. Other works have combined SSL and WSL to learn more generalized representations [37, 17, 53]. Xin Yuan et al. [53] combine vision-language training with contrastive SSL to learn visual representations from multimodal, image-text data. SLIP [37] further builds on this work by showing that combining SimCLR [9] with CLIP [40] objectives outperforms both method in isolation by relatively large margins across multiple downstream tasks. In addition, MaskCLIP [17] presents a masked self-distillation approach combined with contrastive vision-language pre-training, further outperforming CLIP and other joint WSL and SSL methods.
@@ 3 Harmony
We introduce Harmony, a joint self-supervised and weakly-supervised framework for learning semantic and localized visual representations in the wild. The main components of our framework are shown in Figure 1. Namely we define vision student and teacher encoders, ${E}_{V}$ and $\bar{E}_{V}$, text student and teacher encoders, ${E}_{T}$ and $\bar{E}_{T}$, and vision and text decoders, $D_{V}$ and $D_{T}$. Our teacher encoders are used to generate self-distillation and soft targets, and are defined as being the exponential moving average (EMA) of the students. The vision and text decoders are used to map from embedding space to pixel and word token space, respectively. All components use the Transformer [46] architecture, and we use the standard vision transformer (ViT) implementation for processing images [18].
Our framework optimizes five different objectives simultaneously with the goal of learning robust general-purpose visual representations from web-scraped images-caption pairs. Our losses are (1) text-guided contrastive learning that is identical to CLIP [40] but with added soft targets, (2) feature self-distillation following iBOT [55], (3) pixel prediction following MAE [24], word prediction or MLM [15], and (5) text self-distillation, which is similar to iBOT’s patch-level objective but applied on word embeddings [55]. In the following sections, we detail each objective in our framework.
@@ 3.1 Text-guided Contrastive Learning with Soft Targets
We maximize the similarity between paired embedding sets, $\mathbf{v}_{i}$ and $\mathbf{t}_{j}$ where $i=j$ and minimize the similarity between unpaired sets where $i\neq j$. More formally, we define the InfoNCE loss as our training objective [45] following CLIP [40], where $\mathcal{L}_{\text{InfoNCE}}=\mathcal{L}_{v}+\mathcal{L}_{t}$ and
@@
$P_{v}(v_{i},t_{j};\tau)=\frac{\exp(\text{sim}(\mathbf{v_{i}},\mathbf{t_{j}})/\tau)}{\sum_{k=1}^{N}\exp(\text{sim}(\mathbf{v_{i}},\mathbf{t_{k}})/\tau)}\,.$ (2)
The function $\text{sim}(\mathbf{v_{i},t_{j}})$ is the cosine similarity, $\text{sim}(\mathbf{v_{i},t_{j}})=\mathbf{v_{i}}^{T}\mathbf{t_{j}}$, and $\tau$ is a learnable temperature parameter. The loss $\mathcal{L}_{t}$ and function $P_{t}$ are defined in a symmetrical way [45]. Since the above $\mathcal{L}_{\text{InfoNCE}}$ uses hard targets (1s and 0s), we will refer to it as $\mathcal{L}_{\text{Hard}}$.
Notice that $\mathcal{L}_{v}$ (and $\mathcal{L}_{t}$ given the symmetry) is the cross-entropy function $\mathcal{H}(a,b)=-a\log b$ applied across images caption pairs. Because of that, we can rewrite $\mathcal{L}_{v}$ as $\mathcal{H}(\textbf{I}_{\textbf{N}},P(\textbf{V}\textbf{T}^{T};\tau))$ where $\mathbf{V,T}\in\mathbb{R}^{{N\times d}}$ are matrices that contain a batch of image and text embeddings, with $d$ being the embedding size. Therefore, $\mathcal{L}_{\text{Hard}}$ can be rewritten as
@@
Soft targets. The loss function in the original CLIP implementation [40] assumes that there is a one-to-one correspondence between image and caption pairs. This is because the target in the cross entropy function in Equation 3 is the identity matrix $\textbf{I}_{\textbf{N}}$. This means that, given an image embedding $\mathbf{v}_{i}$ and a set of textual embeddings $\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{N}\}$, minimizing the objective can be viewed as an $N$-way multi-class classification problem. This assumption does not always accurately represent the relationship between image-captions sets, especially in the case of uncurated data. Certain captions can describe many different images with varying degrees, regardless of their original pairing, and the target should reflect some function of how semantically similar an image-caption set is.
Recent works have tried to mitigate this issue by incorporating soft instead of hard targets [1, 19, 42]. This is a non-trivial task since generating soft targets assumes some pre-existing knowledge about how semantically similar image-caption pairs are. Gao et al. [19] relies on a pre-trained object detection model to generate soft similarity targets, while [1] used a self-distillation approach to dynamically generate soft targets without pre-training. In the latter approach the same model is used to both generate soft targets and make predictions for calculating the loss, by splitting the mini-batch into student and teacher targets. Instead, we use an EMA self-distillation method for generating soft-targets, motivated by the fact that our framework already defines an EMA model for feature self-distillation in Section 3.2, so the addition of soft targets comes at little computational expense.
To generate soft targets, we define vision and text teacher models, $\bar{E}_{V}$ and $\bar{E}_{T}$, respectively, and their corresponding $\bar{g_{v}}$ and $\bar{g_{t}}$, which are the encoders with the attached projection heads. We do not propagate gradients through the parameters of either $\bar{E}_{V}$ or $\bar{E}_{T}$ and instead update their weights using the EMA of the students. Given the same image-text pair collections $\{(v_{1},t_{1}),(v_{2},t_{2}),...,(v_{N},t_{N})\}$, we generate embedding targets $\mathbf{\bar{v}}_{i}=\bar{g}_{v}(v_{i})$ and $\mathbf{\bar{t}}_{i}=\bar{g_{t}}(t_{i})$. We can represent these targets as the matrices $\bar{\textbf{V}}$ and $\bar{\textbf{T}}$ like in Equation 3. Our soft CLIP loss, $\mathcal{L}_{\text{Soft}}$ is then defined as
$\mathcal{L}_{\text{Soft}}=\mathcal{H}(\bar{\textbf{V}}\bar{\textbf{T}}^{T},P(\textbf{V}\textbf{T}^{T};\bar{\tau}))+\mathcal{H}(\bar{\textbf{T}}\bar{\textbf{V}}^{T},P(\textbf{T}\textbf{V}^{T};\bar{\tau}))\,.$ (4)
@@
$\mathcal{L}_{Contrastive}=$ $\,\alpha_{c}\left[\mathcal{H}(\mathbf{I}_{\mathbf{N}},P(\mathbf{V}\mathbf{T}^{T};\tau))+\mathcal{H}(\mathbf{I}_{\mathbf{N}},P(\mathbf{T}\mathbf{V}^{T};\tau))\right]$
$+(1-\alpha_{c})\left[\mathcal{H}(\bar{\mathbf{V}}\bar{\mathbf{T}}^{T},P(\mathbf{V}\mathbf{T}^{T};\bar{\tau}))+\mathcal{H}(\bar{\mathbf{T}}\bar{\mathbf{V}}^{T},P(\mathbf{T}\mathbf{V}^{T};\bar{\tau}))\right]\,.$ (5)

### v3

@@ Abstract
Vision-language contrastive learning frameworks such as CLIP enable learning representations from natural language supervision and provide strong zero-shot classification capabilities. However, due to the nature of the supervisory signal in these paradigms, they lack the ability to learn localized features, leading to degraded performance on dense prediction tasks such as segmentation and detection. On the other hand, self-supervised learning methods have shown the ability to learn granular representations, complementing the high-level features in vision-language training. In this work, we present Harmony, a framework that combines vision-language training with discriminative and generative self-supervision to learn visual features that can be generalized across different downstream vision tasks. Our framework is specifically designed to work on web-scraped data by not relying on negative examples in the self-supervised learning path and addressing the one-to-one correspondence issue using soft CLIP targets generated by an EMA model. Moreover, Harmony optimizes for five different objectives simultaneously, efficiently utilizing the supervision in each data example, making it even more suited in data-constrained settings. We comprehensively evaluate Harmony across various vision downstream tasks and find that it significantly outperforms the baseline CLIP and outperforms the previously leading joint self- and weakly supervised methods, SLIP, MaskCLIP, and DetailCLIP. Specifically, when compared against these methods, Harmony shows superior performance in linear-probing, fine-tuning, and zero-shot classification on ImageNet-1k, semantic segmentation on ADE20K, and both object detection and instance segmentation on MS-COCO, when pre-training a ViT-B on CC3M. We also show that Harmony outperforms SILC on detection, linear and fine-tuning classification, and outperforms other self-supervised learning methods like iBOT and MAE across all tasks evaluated. Our code is publicly available at https://github.com/MohammedSB/Harmony.
@@ 1 Introduction
Self-supervised and weakly-supervised pre-training have recently shown remarkable success at learning visual representations without direct supervision (Radford et al., 2021; Oquab et al., 2024; Caron et al., 2021b; Chen et al., 2020a; b; He et al., 2020; Grill et al., 2020; Zhou et al., 2022; He et al., 2021; Mu et al., 2021; Dong et al., 2023). As vision training datasets continue to scale, it becomes progressively more difficult and expensive to provide manual supervision in the form of labels, making the development of robust self-supervised learning (SSL) and weakly-supervised learning (WSL) techniques more integral.
Figure 1: Performance comparison across different tasks. Harmony significantly outperforms CLIP, SLIP, MaskCLIP, and DetailCLIP across all tasks, and generally has a larger improvement in performance as we scale the ViT size, except for DetailCLIP which is low performing at ViT-S. Moreover, for classification and detection, the small version of Harmony outperforms the base version of the other methods. We used MS-COCO for object detection, ADE20K for semantic segmentation, and ImageNet for classification, using linear-probing. See Table 1 for numerical values.
@@
Weakly-supervised learning, specifically language-guided or language-supervised learning, was popularized by CLIP (Radford et al., 2021) and learns visual and textual representations using contrastive loss by maximizing the similarity between image-captions pairs and minimizing the similarity of non-paired image-captions (Radford et al., 2021; Cherti et al., 2022). Because this approach relies on semantic captions as a supervisory signal, language-supervised models are strong at high-level tasks like image classification, but significantly underperform on dense, low-level prediction tasks that require localized features (Radford et al., 2021; Wang et al., 2024; 2022). In other words, these paradigms are good at learning what objects are present in a visual input, but not where they are. One approach of introducing local information into WSL frameworks is to combine it with self-supervised learning (Dong et al., 2023; Mu et al., 2021; Yuan et al., 2021). Unlike language-supervised learning that maps across modalities (e.g. image to text), SSL maps to the same visual modality, making it more granular and localized for visual tasks (Caron et al., 2021b; He et al., 2021; Assran et al., 2023).
@@
Recent works in self-supervised learning formulate the pre-training task as either discriminative or generative (Ozbulak et al., 2023; Doersch et al., 2016). For discriminative SSL methods, the model learns visual representations from images by differentiating between positive images pairs, and optionally repelling negative pairs (Chen et al., 2020a; Grill et al., 2020; Caron et al., 2021b; Oquab et al., 2024; He et al., 2020; Chen et al., 2020b). On the other hand, generative approaches learn visual representations by masking certain parts of an image and learning to reconstruct the missing parts, given the original image (He et al., 2021; Tong et al., 2022; Gupta et al., 2023). Because SSL methods do not require manually annotated labels, they can be used for training large neural networks on huge image datasets scraped from the web (Oquab et al., 2024).
@@
Combining language-supervised and self-supervised learning has led to superior learning paradigms (Dong et al., 2023; Mu et al., 2021; Yuan et al., 2021; Naeem et al., 2024; Monsefi et al., 2024). For example, SLIP (Mu et al., 2021) combines CLIP with SimCLR, leading to better zero-shot capabilities and overall more accurate models across multiple downstream tasks. MaskCLIP (Dong et al., 2023) combines CLIP with masked self-distillation, leading to further improvements in a variety of downstream evaluations. More recently, SILC (Naeem et al., 2024) combined CLIP with DINO (Caron et al., 2021b) to learn more localized image features for dense prediction tasks. DetailCLIP (Monsefi et al., 2024) combines A-CLIP (Yang et al., 2023) with MAE (He et al., 2021) and iBOT (Zhou et al., 2022), outperforming all three methods in isolation.
@@
Building on the aforementioned works, we present a novel framework, which we will call Harmony, that combines discriminative and generative self-supervision with language-supervised learning in order to learn general purpose visual representations from image-captions pairs gathered from the web. Our approach is designed to learn semantic visual representations useful for high-level vision tasks and fine-grained, low-level representation for dense prediction tasks at the same time.
@@
Previous approaches like MaskCLIP (Dong et al., 2023) and SLIP (Mu et al., 2021) rely on one-to-one correspondences and hard negative examples which fail to model the inherent semantic relationship between non-paired samples (Andonian et al., 2022). This is because for a certain image, captions of other images in the batch could still describe this image with varying degrees, especially when the batch size is large. Moreover, particularly in an uncurated data setting, the caption paired with a certain image could simply be incorrect, or the description could be only loosely associated with the image. To remedy these issues, we designed our framework to utilize one-to-many relationships by incorporating soft CLIP targets, and to not rely on negative examples by using self-distillation methods like iBOT (Zhou et al., 2022) rather than SimCLR (Chen et al., 2020a) like in SLIP (Mu et al., 2021). We argue that this makes Harmony better designed for web-scraped data like CC3M (Sharma et al., 2018).
Figure 2: Harmony compared to previous methods. We show our Harmony approach next to previously leading joint methods MaskCLIP (Dong et al., 2023) and SLIP (Mu et al., 2021). Harmony optimizes five different objectives simultaneously, outperforming the other approaches across all vision downstream tasks.
@@
  • We present Harmony, a joint self-supervised and weakly-supervised framework that learns global and local features, generalizing across vision tasks including classification, segmentation, and detection. Harmony optimizes for five different objectives simultaneously, fully leveraging each data example, which makes our method particularly suited for data-constrained settings.
@@
  • Harmony is built to avoid negative examples in the self-supervised learning path and reduce reliance on strict one-to-one correspondences in the text-guided contrastive path, achieving superior or competitive performance with leading methods when pre-trained on the web-scraped CC3M dataset.
  • We provide a unified codebase for all methods that is composable, allowing users to combine any subset of our predefined set of methods, and with our own implementation of SILC and MaskCLIP, which were not previously open-sourced.
@@ 2 Related Works
Discriminative self-supervision. Recent discriminative SSL approaches learn visual representations by discriminating between images that are positive pairs of each other, usually defined as different augmentations of the same image, and optionally pushing away negative pairs. (Grill et al., 2020; Caron et al., 2021b; He et al., 2020; Chen et al., 2020b; a). Methods like SimCLR (Chen et al., 2020a) and MoCo (He et al., 2020; Chen et al., 2020b) use contrastive learning to maximize the embedding similarity between positive pairs while simultaneously reducing the similarity of negative pairs. These approaches ignore the inherent similarity between different images and treat the discriminative problem as a one-to-one scheme, where an image is only similar to its positive pair. Other approaches like BYOL (Grill et al., 2020) and DINO (Caron et al., 2021b) use self-distillation to only maximize agreement between positive pairs, avoiding that issue. These methods use a momentum encoder, usually refereed to as a target or teacher encoder, to generate embedding targets for an online or student encoder that is trained simultaneously with the teacher encoder. The learning task in these paradigms is to maximize the similarity between embeddings from both the teacher and student encoders given positive image pairs. Self-distillation methods regularly outperform contrastive methods, and reach accuracies that are competitive with supervised learning (Caron et al., 2021b; Zhou et al., 2022; Grill et al., 2020).
@@
Generative self-supervision. Rather than discriminating between images, generative SSL methods learn visual representations by masking certain parts of an images (or certain images in a video) and then learn to reconstruct the masked portions given the original signal (He et al., 2021; Tong et al., 2022; Gupta et al., 2023). This reconstruction task can be viewed as a proxy for learning about visual features from images, improving representation ability and, therefore, performance on downstream tasks. A common approach in this paradigm is called Masked Autoencoder (MAE) (He et al., 2021), which utilizes ViT’s patching paradigm to mask a certain percentage of patches (usually 75%) and minimizes the L2 loss on predicted pixels from the masked patches. Generative SSL approaches are highly scalable due to their lightweight memory usage compared to discriminative approaches, and reach fine-tuning accuracies rivaling those methods (He et al., 2021).
@@
Joint self-supervision methods. Many previous works have combined discriminative and generative self-supervised learning objectives, resulting in improved performance on downstream classification and semantic segmentation tasks (Chen et al., 2023; Huang et al., 2024; Mishra et al., 2022). CMAE (Huang et al., 2024) and CAE (Chen et al., 2023) combine MAE’s pixel reconstruction task with contrastive learning, outperforming either method in isolation. CAN (Mishra et al., 2022) further add a noise prediction task and masks both views in the contrastive learning objective. These works highlight the fact that SSL paradigms can be complementary to each other.
@@
Language-guided and self-supervised learning. Other works have combined SSL and language-guided learning to learn more generalized representations (Mu et al., 2021; Dong et al., 2023; Yuan et al., 2021). Yuan et al. (2021) combine vision-language training with contrastive SSL to learn visual representations from multimodal, image-text data. SLIP (Mu et al., 2021) builds on this work further by showing that combining SimCLR (Chen et al., 2020a) with CLIP (Radford et al., 2021) objectives outperforms both methods in isolation by relatively large margins across multiple downstream tasks. SILC (Naeem et al., 2024) replaces SimCLR with DINO in SLIP, outperforming SLIP. In addition, MaskCLIP (Dong et al., 2023) presents a masked self-distillation approach combined with contrastive vision-language pre-training, further outperforming CLIP and other joint WSL and SSL methods. More recently, DetailCLIP (Monsefi et al., 2024) combines A-CLIP (Yang et al., 2023) with MAE and iBOT to learn more localized visual features for fine-grained tasks. We further build on these works by going an extra step of combining five different objectives, including self-distillation, pixel reconstruction, text-guided contrastive learning, and two additional text-only losses: masked-language modeling and text self-distillation.
@@ 3 Harmony
We introduce Harmony, a joint self-supervised and weakly-supervised framework for learning semantic and localized visual representations in the wild. The main components of our framework are shown in Figure 2. Namely we define vision student and teacher encoders, ${E}_{V}$ and $\bar{E}_{V}$, text student and teacher encoders, ${E}_{T}$ and $\bar{E}_{T}$, and vision and text decoders, $D_{V}$ and $D_{T}$. Our teacher encoders are used to generate self-distillation and soft targets, and are defined as being the exponential moving average (EMA) of the students. The vision and text decoders are used to map from embedding space to pixel and word token space, respectively. All components use the Transformer (Vaswani et al., 2023) architecture, and we use the standard vision transformer (ViT) implementation for processing images (Dosovitskiy et al., 2021).
Our framework optimizes five different objectives simultaneously with the goal of learning robust general-purpose visual representations from web-scraped images-caption pairs. Our losses are (1) text-guided contrastive learning that is identical to CLIP (Radford et al., 2021) but with added soft targets, (2) feature self-distillation following iBOT (Zhou et al., 2022), (3) pixel prediction following MAE (He et al., 2021), (4) word prediction or MLM (Devlin et al., 2019), and (5) text self-distillation, which is similar to iBOT’s patch-level objective but applied on word embeddings (Zhou et al., 2022). In the following sections, we detail each objective in our framework.
@@ 3.1 Text-guided Contrastive Learning with Soft Targets
We maximize the similarity between paired embedding sets, $\mathbf{v}_{i}$ and $\mathbf{t}_{j}$ where $i=j$ and minimize the similarity between unpaired sets where $i\neq j$. More formally, we define the InfoNCE loss as our training objective van den Oord et al. (2019) following CLIP Radford et al. (2021), where $\mathcal{L}_{\text{InfoNCE}}=\mathcal{L}_{v}+\mathcal{L}_{t}$ and
@@
$P_{v}(v_{i},t_{j};\tau)=\frac{\exp(\text{cos}(\mathbf{v_{i}},\mathbf{t_{j}})/\tau)}{\sum_{k=1}^{N}\exp(\text{cos}(\mathbf{v_{i}},\mathbf{t_{k}})/\tau)}\,.$ (2)
The function $\text{sim}(\mathbf{v_{i},t_{j}})$ is the cosine similarity, $\text{sim}(\mathbf{v_{i},t_{j}})=\mathbf{v_{i}}^{T}\mathbf{t_{j}}$, and $\tau$ is a learnable temperature parameter. The loss $\mathcal{L}_{t}$ and function $P_{t}$ are defined in a symmetrical way van den Oord et al. (2019). Since the above $\mathcal{L}_{\text{InfoNCE}}$ uses hard targets (1s and 0s), we will refer to it as $\mathcal{L}_{\text{Hard}}$.
Notice that $\mathcal{L}_{v}$ (and $\mathcal{L}_{t}$ given the symmetry) is the cross-entropy function $\mathcal{H}(a,b)=-a\log b$ applied across image caption pairs. Because of that, we can rewrite $\mathcal{L}_{v}$ as $\mathcal{H}(\textbf{I}_{\textbf{N}},P(\textbf{V}\textbf{T}^{T};\tau))$ where $\mathbf{V,T}\in\mathbb{R}^{{N\times d}}$ are matrices that contain a batch of image and text embeddings, with $d$ being the embedding size. $P$ is the vectorized version of $P_{v}$ and $P_{t}$ in Equation 2, where the cosine similarities, $\text{cos}(\mathbf{v_{i}},\mathbf{t_{j}})$, are calcualted through matrix multiplication. Therefore, $\mathcal{L}_{\text{Hard}}$ can be rewritten as
@@
Soft targets. The loss function in the original CLIP implementation Radford et al. (2021) assumes that there is a one-to-one correspondence between image and caption pairs. This is because the target in the cross entropy function in Equation 3 is the identity matrix $\textbf{I}_{\textbf{N}}$. This means that, given an image embedding $\mathbf{v}_{i}$ and a set of textual embeddings $\{\mathbf{t}_{1},\mathbf{t}_{2},\dots,\mathbf{t}_{N}\}$, minimizing the objective can be viewed as an $N$-way multi-class classification problem. This assumption does not always accurately represent the relationship between image-captions sets, especially in the case of uncurated data. Certain captions can describe many different images with varying degrees, regardless of their original pairing, and the target should reflect some function of how semantically similar an image-caption set is.
Recent works have tried to mitigate this issue by incorporating soft instead of hard targets (Andonian et al., 2022; Gao et al., 2023; Scotti et al., 2023). This is a non-trivial task since generating soft targets assumes some pre-existing knowledge about how semantically similar image-caption pairs are. Gao et al. (2023) relies on a pre-trained object detection model to generate soft similarity targets, while Andonian et al. (2022) used a self-distillation approach to dynamically generate soft targets without pre-training. In the latter approach the same model is used to both generate soft targets and make predictions for calculating the loss, by splitting the mini-batch into student and teacher targets. Instead, we use an EMA self-distillation method for generating soft-targets, motivated by the fact that our framework already defines an EMA model for feature self-distillation in Section 3.2, so the addition of soft targets comes at little computational expense.
To generate soft targets, we define vision and text teacher models, $\bar{E}_{V}$ and $\bar{E}_{T}$, respectively, and their corresponding $\bar{g_{v}}$ and $\bar{g_{t}}$, which are the encoders with the attached projection heads. We do not propagate gradients through the parameters of either $\bar{E}_{V}$ or $\bar{E}_{T}$ and instead update their weights using the EMA of the students. Given the same image-text pair collections $\{(v_{1},t_{1}),(v_{2},t_{2}),...,(v_{N},t_{N})\}$, we generate embedding targets $\mathbf{\bar{v}}_{i}=\bar{g}_{v}(v_{i})$ and $\mathbf{\bar{t}}_{i}=\bar{g_{t}}(t_{i})$. We can represent these targets as the matrices $\bar{\textbf{V}}$ and $\bar{\textbf{T}}$ like in Equation 3, with $\textbf{A}_{V}=P(\bar{\textbf{V}}\bar{\textbf{T}}^{T};\bar{\tau})$ and $\textbf{A}_{T}=P(\bar{\textbf{T}}\bar{\textbf{V}}^{T};\bar{\tau})$. Our soft CLIP loss, $\mathcal{L}_{\text{Soft}}$ is then defined as
$\mathcal{L}_{\text{Soft}}=\mathcal{H}(\textbf{A}_{V},P(\textbf{V}\textbf{T}^{T};\tau))+\mathcal{H}(\textbf{A}_{T},P(\textbf{T}\textbf{V}^{T};\tau))\,.$ (4)
@@
$\mathcal{L}_{C}=\,\alpha_{c}\mathcal{L}_{\text{Hard}}+(1-\alpha_{c})\mathcal{L}_{\text{Soft}}.$ (5)